OpenStudy (anonymous):

The magnitudes of the electric potential and if the electrostatic force depend on which factors

OpenStudy (anonymous):

A. Mass and distance B. Distance and amount of charge C. Mass and amount of charge D. Density and amount of charge

OpenStudy (anonymous):

a, c, d) is it gravity, or electricity? The SOURCE for everything electrical is electric charge. ... too bad they ignored the electric permitivity or k (absolutely necessary!)

OpenStudy (radar):

I think it is charge, that is why B needs further consideration.

OpenStudy (anonymous):

and Force comes from closeness (proximity), not distance.

OpenStudy (irishboy123):

\(\vec F = k \frac{Q_1 \ Q_2}{r^2} \hat r\) \(V = -\frac{k \ Q}{r}\)

OpenStudy (ybarrap):

@IrishBoay123 's has got it, charge and distance are the variables that determine electrostatic force and electric potential. See my favorite lecture on the subject - Lec 04: Electrostatic Potential and Electric Energy | 8.02 Electricity and Magnetism (Walter Lewin) https://youtu.be/5aQdZ26uR7E